OpenStudy (anonymous):

Show that (5 + 3√2)^2 = 43 + 30√2.

OpenStudy (anonymous):

foil the left side

OpenStudy (anonymous):

exapnd lhs using (a+b)^2=a^2+b^2 +2ab

OpenStudy (jiteshmeghwal9):

No don't FOIL, it will take a long time so, use the identity (a+b)^2=a^2+b^2+2ab on the left side (5)^2+(3sqrt{2})^2+2(5)(3sqrt{2})=43+30sqrt{2}

OpenStudy (jiteshmeghwal9):

Now solve the left side u will probably gt R.H.S.=L.H.S.

OpenStudy (anonymous):

that's exactly what foil is though, expanding it

OpenStudy (whpalmer4):

and it's easily done in this case! 5*5 + 5*3*sqrt(2) + 5*3*sqrt(2) + 3*3*sqrt(2)*sqrt(2) = 25 + 30*sqrt(2) + 18 = 43 + 30*sqrt(2)

OpenStudy (anonymous):

Thank you so much! :)